ProForm acquired 80 percent or ClipRite on June 30, 2017, for $1,280,000 in cash. Based on ClipRites acquisition-date fair value, an unrecorded intangible of $800,000 was recognized and lis being amortized at the rate of $14,000 per year. No goodwill was recognized in the acquisition. The noncontrolling interest fair value was assessed at $320,000 at the acquisition date. The 2018 financial statements are as Tollows: ProForm clipRite Sales (980,000) 625,000 280,000 (960,000) 490,000 190,000 operating expenses Dividend income 80,000 Net income Retained earnings Net income 155,000) (280,000) $(2,500,000) 1,030,000) (280,000) $(2,375,000) 1,210,000) 155,000) 280,000 Dividends declared 100,000 Retained earnings, 12/31/18 Cash and receivables 580,000 480,000 Inventory Investment in ClipRite Fixed assets Accumulated depreciation 470,000 ,280,000 1,500,000 890,000 1,500,000 (400,000) (450,000) $ 3,430,000 2,410,000 Totals Liabilities Common stock Retained earnings 12/31/18 (555,000) (500,000) (700,000) (500,000) (2,375,000) (1,210,000) $3,430,000) (2,410,000) Totals ProForm sold ClipRite inventory costing $87000 during the last six months of 2017 for $270,000. At year-end, 30 percent remained. ProForm sells CipRite inventory costing $290,000 during 2018 for $430,000. At year-end, 10 percent is left. Determine the consolidated balances for the following accounts: Sales Cost of goods sold 0 Dividend income Net income attributable to noncontrolling interest Inven 1.510.000 644,100 484,000 interest in subsidiary, 12/3118
Answer #1

Intra-Entity gross profit, 12/31/17: (upstream transfer)

Intra-entity gross profit ($270,000 – $87,000)

183000

Inventory remaining at year's end

20%

Intra-entity gross profit in inventory, 12/31/17

36600

Intra-Entity gross profit, 12/31/18: (upstream transfer)

Intra-entity gross profit ($430,000 – $290,000)

140000

Inventory remaining at year's end

10%

Intra-entity gross profit in inventory, 12/31/18

14000

Sales = Proform's Sales + ClipRite's Sales - Intra-entity transfer = $980,000 + $960,000 - $430,000 = $1,510,000

Cost of goods sold = $662400

Proform's COGS book value

625000

ClipRite's COGS book value

490000

Eliminate intra-entity transfers

(430000)

Recognized gross profit deferred in 2017

(36600)

Deferral of 2018 Intra-entity gross profit in inventory

14000

Consolidated cost of goods sold

662400

Operating expenses = Proform's Operating Expenses + ClipRite's Operating Expenses + Intangible Amortization = 280,000+190,000+14,000 = $484,000

Dividend income = 0 (intra-entity transfer eliminated in consolidation)

Net income attributable to noncontrolling interest:

ClipRite reported net income for 2018

280000

Intangible amortization

(14000)

2017 gross profit recognized in 2018

36600

2018 gross profit deferred

(14000)

ClipRite adjusted net income for 2018

288600

Outside ownership

20%

Net income attributable to noncontrolling interest

57720

Inventory = Proform's Inventory + ClipRite's Inventory - Ending intra-entity gross profit = 470000 +880000 -14000 = $1336000

Noncontrolling interest in subsidiary, 12/31/18

20% beginning book value less $36600 intra-entity gross profit ((2410000-280000)-36600)*20%)

418680

Excess intangible allocation (20% × (800000-14000))

157200

Net income attributable to noncontrolling interest

57720

Dividends (20% × $100,000)

(20000)

Total noncontrolling interest at 12/31/18

613600

Therefore,

Sales

$1,510,000

Cost of goods sold

$662400

Operating expenses

$484,000

Dividend income

$0

Net income attributable to noncontrolling interest

$57720

Inventory

$1336000

Noncontrolling interest at 12/31/18

$613600